Mean of the sampling distribution of the sample mean


1
Expert's answer
2021-05-07T10:14:10-0400

The Sampling Distribution of the Sample Mean is the distribution of all possible sample means of a given sample size.

For Example:

A population consists of the numbers 2, 4, 8, 10 and 5

Sampling distribution of size 3 can be given as




"P(\bar X=4.66) = \dfrac{1}{5}"


"P(\bar X=6.66) = \dfrac{1}{10}"


"P(\bar X=5.66) = \dfrac{1}{10}"


"P(\bar X=5) = \dfrac{1}{5}"


"P(\bar X=7.33) = \dfrac{1}{10}"


"P(\bar X=6.33) = \dfrac{1}{10}"


"P(\bar X=3.66) = \dfrac{1}{10}"


"P(\bar X=7.66) = \dfrac{1}{10}"